OpenStudy (anonymous):

1.Factorise using the remainder theorem 2x^3 - 9x^2 - 11x + 30. 2.Factorise : (2y+x)^2(y-2x)+(2x+y)^2(2x-y)

OpenStudy (anonymous):

plzz answer..>!!:((

OpenStudy (anonymous):

@jdoe0001 plzz cn u answer any 1.?? plzz

OpenStudy (jdoe0001):

hmmm, do you know what the remainder theorem is? because if the exercises asks for it,... it's assume you're supposed to know it already :S

OpenStudy (anonymous):

ik wat is remainder theoram..bt i dk hw to do dis 1.. HELP:( ;9

OpenStudy (jdoe0001):

well,... I looked and looked, and I don't think the remainder ALONE will do it the way I've do it is using the "rational root test" and the rational root test uses the remainder theorem but just a bit

OpenStudy (anonymous):

k..cn u shw hw??

OpenStudy (anonymous):

@jdoe0001 cn u help me??

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=GTx-rXTQg1o <---

OpenStudy (jdoe0001):

check that, let me know for anything confusing :)

OpenStudy (anonymous):

do u noe hw to do d second question?

OpenStudy (jdoe0001):

well, the second is just factoring you have 2 polynomials being + ADDED, so the way to factor them is to get "common factor" on both

OpenStudy (anonymous):

cn u shw hw??plzz step by step

OpenStudy (jdoe0001):

ok, lemme give you.... a quick example ab + bc what would be the common factor there? :)

OpenStudy (anonymous):

b

OpenStudy (jdoe0001):

yeah, so lemme show you the one you have there now \(\bf \color{blue}{(2y+x)^2}(y-2x)+\color{blue}{(2x+y)^2}(2x-y)\) see any common factors?

OpenStudy (anonymous):

2

OpenStudy (jdoe0001):

so, you take out the common factors out

OpenStudy (anonymous):

k den

OpenStudy (jdoe0001):

then, that's about it on factoring, you end up with several "factors" or item multiplying each other

OpenStudy (anonymous):

wat do u gt?

OpenStudy (anonymous):

idk hw to do it..cn u shw fully hw to do dis equation??

OpenStudy (jdoe0001):

in the previous example ab + bc # common factor is "b" so if we take out "b", that is, we take away "b" from "ab" and leave only "a" by its lonesome and take away "b" from "bc" and leave "c" all by its lonesome too we're left with \(\bf b(a+c)\)

OpenStudy (jdoe0001):

so you'd do the same with any polynomial to get a common factor

OpenStudy (jdoe0001):

some are just longer than just 2 terms, it could be any amount of terms

OpenStudy (anonymous):

ya...i gt dat..bt hw do u do ni dis equation is it 2[(y+x)(y-x)+(x+y)(x-y)] =2(y+x+y) =2y^2x

OpenStudy (anonymous):

u cn do cancellation..rie??

OpenStudy (jdoe0001):

\(\bf (2y+x)^2(y-2x)+(2x+y)^2(2x-y)\\ \textit{let's call }(2y+x)^2 = a\\ a(y-2x)+a(2x-y) \implies a[(y-2x)+(2x-y) ]\\ \implies (2y+x)^2[(y-2x)+(2x-y) ]\)

OpenStudy (anonymous):

ohhk

OpenStudy (anonymous):

so d final anser is dat??

OpenStudy (jdoe0001):

on another note, if you were to expand say -(2x -y) what would it look like?

OpenStudy (anonymous):

i hv another questin

OpenStudy (anonymous):



OpenStudy (jdoe0001):

can't say right off, might be better to post anew so we all can see and help and revise each other

OpenStudy (anonymous):

hmmk